Problem 3 A parallel-plate capacitor with plate area 2.20 cm2 and air-gap separation 0.13 mm is connected to a 12.00 V battery, and fully charged. The battery is then disconnectecd (a) What is the charge on the capacitor? 1.80x10-10 c You are correct. Computers answer now shown above. Your receipt is 160-5206 Previous Tries (b) The plates are now pulled to a separation of 0.37 mm. What is the charge on the capacitor now? 180x1010 c You are correct. Computers answer now shown above Your receipt is 160-8526 (c) What is the potential difference across the plates now? 3.42x101v You are correct. Computers answer now shown above. Your receipt is i 60-6167 Previous Tries (d) How much work was required to pull the plates to their new separation? 6.156 10-9 J Submit Answer Incorrect. Tries 2/6 Previous Tries
Answer #1

d)

W = final energy - initial energy

W = 0.5* QV - 0.5* qv

W = 0.5* 1.8*10^-10* 34.2 - 0.5* 1.8* 10^-10* 12

W = 1.998*10^-9 J

=========